what is x^4/3 - 13x^2/3 + 36 = 0
Answer by Alan3354(69443)

what is x^4/3 - 13x^2/3 + 36 = 0
----------
It's a quadratic equation in x^(2/3).
If you want to solve for x (you didn't say):
Sub another name for x^(2/3), say z.
--> z^2 - 13z + 36 = 0
(z-4)*(z-9) = 0
z = 4, 9
--------
Now sub back
x^(2/3) = 4
x^2 = 64
x = +8, -8
-------------
x^(2/3) = 9
x^2 = 729
x = +27, -27